Calculate pressure in atmospheres, required to compress a sample of helium gas from 26.6L (at 1.00atm) to 3.33L at constant temperature
1
Expert's answer
2020-04-18T07:07:23-0400

From gas law p*V = n*R*T = constant at constant temperature.

p1*V1 = p2*V2

Pressure needed to compress 26.6 L (1 atm) to 3.33 L is:

p = 1*26.6/3.33 = 7.99 atm